describe what steps you would take to solve 8x+5=2x-7
Answer by likaaka(51)

Ultimately you are trying to find the value of the variable x, so you need to isolate the x meaning put all x's on one side of the equal sign and all known values on the other side of the equal sign.
So lets put all x's on the right by subtracting the 2x on the left from BOTH sides of the equation
8x+5-2x=2x-7-2x
Simplify by combining like terms 8x-2x is 6x on the right and 2x-2x is 0 on the left and the resulting equations becomes
6x+5=-7
Now you put the known values on the left by subtracting the 5 on the right from BOTH sides of the equation
6x+5-5=-7-5
Simplify by combining like terms 5-5 is 0 on the right and -7-5 is -12 on the left and the resulting equation becomes
6x=-12
Then we solve for x by dividing BOTH sides of the equation by 6
(6x)/6=12/6
x=2
In order for equation to remain true, what you do to one side you must do the the other side of the equal sign